# Sensing with submarine optical cables

### v2

@@ II Basic equations
representing an anticlockwise rotation by the angle $\theta(z)$ of in the $x-y$ plane with respect to the $z$ axis, so that $\mathbf{V}(z)\,\mathbf{B}(z)\,\mathbf{V}^{\dagger}(z)$ produces a rotation by the same angle of the birefringence vector, and
@@ IV Phase-based sensing
Phase sensing relies on measuring the perturbation to the phase of an optical field propagating along an optical fiber due to strain induced by coupling with the surrounding environment. In the previous section, it was demonstrated that strain causes a local alteration in the polarization-averaged wavevector, as described by Equation (33). Consequently, this modification affects the polarization-averaged phase accumulated over a distance $z$ adding to Eq. (24) the term
@@
The primary noise source that limits the sensitivity of phase measurements is the phase noise of the laser probe. In the following, we will examine the limitations that phase noise imposes on the sensitivity of phase measurements. Consider a laser with quantum-limited white frequency noise. Accurate detection of the phase deviation occurring over the time $T$
@@
requires that $\Delta\varphi_{\mathrm{sig}}(T)\gg\Delta\varphi_{\mathrm{rms}}$, where $\Delta\varphi_{\mathrm{rms}}$ is the root-mean square of the phase deviation due to the laser phase noise. The phase fluctuations of a quantum limited laser are described by a Wiener process with diffusion constant $D_{\varphi}=2\pi\nu$, where $\nu$ is the laser linewidth. After a time $T$ the root-mean square phase deviations originating from the laser phase noise are therefore
$\Delta\varphi_{\mathrm{rms}}(T)=\langle\Delta\varphi_{\nu}(T)\rangle^{1/2}=\sqrt{D_{\varphi}T}.$ (41)
The condition $\Delta\varphi_{\mathrm{sig}}(T)\gg\Delta\varphi_{\mathrm{rms}}$ becomes
$\Delta\varphi_{\mathrm{sig}}(T)\gg\sqrt{D_{\varphi}T}.$ (42)
@@
$T\gg\frac{2\pi\nu}{\dot{\varphi}^{2}}.$ (44)
Another condition that should be met for the $2\pi$ periodicity of the phase is that $\Delta\varphi=\dot{\varphi}T\leq 2\pi$, that is $T\leq 2\pi/\dot{\varphi}$. Combination of the two conditions gives
$\frac{2\pi}{\dot{\varphi}}\geq T\gg\frac{2\pi\nu}{\dot{\varphi}^{2}},$ (45)
so that we obtain from the first and last term of the inequality the somehow expected result that
$\nu\ll\dot{\varphi}.$ (46)
This condition was verified in the earthquake detection cases presented in ref. Marra et al., 2018, wherein the earthquakes induced a phase dynamics covering intervals on the order of thousands of radians over a timescale of tens of seconds, and the laser employed had sub-hertz equivalent linewidth. It is questionable whether the detection of processes with much weaker dynamics that occur over the same timescale or longer, like those associated with tsunami’s propagation, would be possible even utilizing lasers with similar ultra-low linewidth.
@@
and where $N_{0}(t)$ is the phase noise of a quantum limited laser with linewidth $\nu$, namely a white noise term with correlation function
@@
We employed a normalization factor $1/n$ in the Fourier transform to ensure that the spectral amplitude $\tilde{\varphi}(\Omega_{\overline{k}})$ of a pure sinusoidal modulation at frequency $\Omega_{\overline{k}}$ is independent of the number of samples $n$. Fourier transforming both sides of Eq. (50) yields
@@
The use of this equation and of Eq. (52) yields
@@
A faithful detection of the phase modulation at angular frequency $\Omega_{k}$ requires that in Eq. (55) the signal is much larger than the noise,
$\langle|\tilde{N}(\Omega_{k})|^{2}\rangle\ll|\Omega_{k}|^{2}|\tilde{\varphi}(\Omega_{k})|^{2},$ (58)
that is, using Eq. (57)
$\nu\ll\frac{1}{2\pi}|\Omega_{k}|^{2}|\tilde{\varphi}(\Omega_{k})|^{2}nT.$ (59)
Equation (54) also dictates that $|\Omega_{k}|$ must not exceed $|\Omega_{n/2}|$, implying
$T\leq\frac{\pi}{|\Omega_{k}|},$ (60)
so that we may replace $T$ with $\pi/|\Omega_{k}|$ at the right-hand side of Eq. (59). This substitution yields the final result
$\nu\ll\frac{n}{2}|\Omega_{k}||\tilde{\varphi}(\Omega_{k})|^{2}.$ (61)
The minimum laser linewidth increases linearly with the signal frequency, the signal spectral power, and the number of samples used in the short-time Fourier transform. Equation (61) illustrates that employing a large number of samples $n$ can partially alleviate the requirements on the laser linewidth, which were more stringent for the detection of transient phase deviations, as shown at the beginning of this section. However, analysis of Eq. (59) reveals that this can only be achieved by widening the time window of the short-time Fourier transform $nT$, thus sacrificing temporal resolution. Consequently, the necessity of using ultra-stable lasers for detecting low intensity sub-hertz signals persists even with approaches based on the short-time Fourier transform.
@@ V Polarization-based sensing
with $\langle\tau\rangle^{2}=8\langle\tau^{2}\rangle/(3\pi)$ the mean polarization mode dispersion square.Galtarossa et al. (2001) If we define as $\kappa^{2}=\langle\tau\rangle^{2}/z$ the averaged square polarization mode dispersion of the fiber in ps/$\sqrt{\mathrm{km}}$, we may eliminate $2L_{\mathrm{f}}\langle\beta^{2}\rangle$ in the correlation functions by using
@@
[…]
@@
[…]
@@ VI Comparison between phase and polarization accumulation
Assuming $\kappa=0.03$ ps/$\sqrt{\mathrm{km}}$ and $n_{0}=1.5$, we obtain for $L_{\mathrm{f}}=2\Lambda_{0}/\cos\vartheta$ and $L_{f}=400$ km, $\Delta\varphi(z,t)\simeq 6000\,\langle|\Delta\vec{s}(z,t)|^{2}\rangle^{1/2}$ rad. When the ratio $L_{\mathrm{f}}/(\Lambda/\cos\vartheta)$ becomes larger, the efficiency of the phase modulation drops very rapidly as a consequence of the Gaussian profile and the fact that we assumed $\Lambda$ as a constant, so that the expression may become meaningless in this limit.
@@ VIII Localization with polarization
Assume that $\mathbf{U}_{\mathrm{f}}$ is the Jones matrix that describes the evolution of the polarization of a single mode fiber up to a given distance $z$ and $\Delta\mathbf{U}$ the unitary matrix describing the evolution in a section that goes from $z$ to $z+\Delta z$ along the same fiber, which we will refer to as the section of interest in the following. Then, the backscattered field is either rerouted through a different fiber with Jones matrix $\mathbf{U}_{\mathrm{b}}$ or transmitted back through the same fiber, in which case $\mathbf{U}_{\mathrm{b}}=\mathbf{U}_{\mathrm{f}}$. The Jones matrix $\mathbf{U}_{\mathrm{rt}}(1)$ describing the round-trip propagation from 00 to $z$ and back, and the Jones matrix $\mathbf{U}_{\mathrm{rt}}(2)$ describing the round-trip propagation from 00 to $z+\Delta z$ and back, areGaltarossa et al. (2008)
@@ IX Environmental sensing using the state of polarization
$\left\langle\left[\mathbf{R}_{\omega_{2}}^{-1}(z^{\prime})\vec{\beta}(z^{\prime})\times\vec{s}_{0}\right]\cdot\left[\mathbf{R}_{\omega_{1}}^{-1}(z^{\prime\prime})\vec{\beta}(z^{\prime\prime})\times\vec{s}_{0}\right]\right\rangle.$
The result of the average is (see appendix A for the detailed derivation)
@@
$\frac{\pi}{4}\left(\frac{2\pi c}{\lambda}\right)^{2}\kappa^{2}\int_{0}^{z}{\operatorname{d}}z^{\prime}\epsilon(z^{\prime},t^{\prime})\epsilon(z^{\prime},t)F(z^{\prime}),$
@@
Expressing $F(z)=\exp(-z/\Delta Z)$ with $\Delta Z=2/(\pi^{3}\Delta f^{2}\kappa^{2})$, and using the values of the systems of ref. Zhan, 2020, $\kappa=0.03$ ps/$\sqrt{\mathrm{km}}$ and $\Delta f=76.5$ GHz, we obtain $\Delta Z\simeq 12,247$ km. In the presence of a localized perturbation centered at $z=z_{p}$ over a spatial extension $\Delta z\ll\Delta Z$, we may replace $F(z^{\prime})$ with $F(z_{p})$ in the integral at right-hand side of Eq. (125), obtaining
@@
$\frac{\pi}{4}\left(\frac{2\pi c}{\lambda}\right)^{2}\xi^{2}\kappa^{2}F(z_{p})\int_{0}^{z}{\operatorname{d}}z^{\prime}\epsilon(z^{\prime},t^{\prime})\epsilon(z^{\prime},t),$
@@
Figure 8: Autocorrelation of $\Delta s_{1}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 552.35 s, from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake 23 June 2020 15:29:05 plotted from $-200$ s to $200$ s (left panel), and a zoom from $-50$ s to $50$ s (right panel).
Figure 9: Autocorrelation of $\Delta s_{2}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 552.35 s, from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake on 23 June 2020, 15:29:05 plotted from $-200$ s to $200$ s (left panel), and a zoom from $-50$ s to $50$ s (right panel).
@@
Figure 16: Autocorrelation of $\Delta s_{1}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 519.9 s (red), from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake, 3 June 2020 07:35:33 plotted from $-200$ s to $200$ s (left panel), and a zoom from $-50$ s to $50$ s (right panel).
Figure 17: Autocorrelation of $\Delta s_{2}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 519.9 s (red), from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake on 3 June 2020, 07:35:33 plotted from $-200$ s to $200$ s (left panel), and a zoom from $-50$ s to $50$ s (right panel).
@@
These observations suggest that, in principle, within a single-pass configuration of an operational transmission system, cross-correlation between two closely spaced channels, whose spacing can be optimized for maximum accuracy, may permit the localization of the position on the link where an earthquake occurs. The optimization involves the choice of a frequency spacing maximizing the sensitivity of $F(z_{p})$ on a $z_{p}$ ranging from 00 to the link span $z$. A good recipe may be setting $\Delta z\simeq z$.
@@ X Conclusions
Appendix A Derivation of Eq. (125)
Equation (124) can be rewritten as

### v3

@@ II Basic equations
representing an anticlockwise rotation by the angle $\theta(z)$ of in the xyx-y plane with respect to the $z$ axis, so that $\mathbf{V}(z)\,\mathbf{B}(z)\,\mathbf{V}^{\dagger}(z)$ produces a rotation by the same angle of the birefringence vector, and
@@ IV Phase-based sensing
Phase sensing relies on measuring the perturbation to the phase of an optical field propagating along an optical fiber due to strain induced by coupling with the surrounding environment. In the previous section, it was demonstrated that strain causes a local alteration in the polarization-averaged wavevector, as described by Eq. (33). Consequently, this modification affects the polarization-averaged phase accumulated over a distance $z$ adding to Eq. (24) the term
@@
The primary noise source that limits the sensitivity of phase measurements is the phase noise of the laser probe. In the following, we examine the limitations that phase noise imposes on the sensitivity of phase measurements. Consider a laser with quantum-limited white frequency noise. Accurate detection of the phase deviation occurring over the time $T$
@@
requires that $|\Delta\varphi_{\mathrm{sig}}(T)|\gg\Delta\varphi_{\mathrm{rms}}$, where $\Delta\varphi_{\mathrm{rms}}$ is the root-mean square of the phase deviation due to the laser phase noise. The phase fluctuations of a quantum limited laser are described by a Wiener process with diffusion constant $D_{\varphi}=2\pi\nu$, where $\nu$ is the laser linewidth. After a time $T$ the root-mean square phase deviations originating from the laser phase noise are therefore
$\Delta\varphi_{\mathrm{rms}}(T)=\langle\Delta\varphi_{\nu}^{2}(T)\rangle^{1/2}=\sqrt{D_{\varphi}T}.$ (41)
The condition $|\Delta\varphi_{\mathrm{sig}}(T)|\gg\Delta\varphi_{\mathrm{rms}}$ becomes
$|\Delta\varphi_{\mathrm{sig}}(T)|\gg\sqrt{D_{\varphi}T}.$ (42)
@@
$\nu\ll\frac{\dot{\varphi}^{2}T}{2\pi}.$ (44)
A longer sampling time effectively averages the effects of laser phase noise, thereby reducing the requirements on the linewidth. However, the $2\pi$ periodicity of the phase imposes that $|\Delta\varphi_{\mathrm{sig}}|=|\dot{\varphi}|T\leq 2\pi$. Using for $T$ its maximum value $2\pi/|\dot{\varphi}|$, we obtain the somehow expected result
$\nu\ll|\dot{\varphi}|.$ (45)
This condition was met in the earthquake detection cases presented in ref. Marra et al., 2018, wherein the earthquakes induced a phase dynamics covering intervals on the order of thousands of radians over a timescale of tens of seconds, and the laser employed had sub-hertz equivalent linewidth. It is questionable whether the detection of processes with much weaker dynamics that occur over the same timescale or longer, like those associated with tsunami’s propagation, would be possible even utilizing lasers with similar ultra-low linewidth.
@@
and where $N_{0}(t)$ is the frequency noise of a quantum limited laser with linewidth $\nu$, namely a white noise term with correlation function
@@
We employed a normalization factor $1/n$ in the Fourier transform definition to ensure that the spectral amplitude $\tilde{\varphi}(\Omega_{\overline{k}})$ of a pure sinusoidal modulation at frequency $\Omega_{\overline{k}}$ is independent of the number of samples $n$. Fourier transforming both sides of Eq. (49) yields
@@
Taking the square of Eq. (51), averaging and using Eq. (55) yields
@@
A faithful detection of the phase modulation at angular frequency $\Omega_{k}$ requires that the signal in Eq. (54) is much larger than the noise,
$\langle|\tilde{N}(\Omega_{k})|^{2}\rangle\ll|\Omega_{k}|^{2}|\tilde{\varphi}(\Omega_{k})|^{2}.$ (57)
By using Eq. (56) for $h=k$ into Eq. (57), solving for $\nu$ and defining $T_{\mathrm{win}}=nT$ as the amplitude of the time window of the Fourier transform and $f_{k}=\Omega_{k}/(2\pi)$ as the frequency in hertz, Eq. (57) yields
$\nu\ll 2\pi|f_{k}|^{2}|\tilde{\varphi}(2\pi f_{k})|^{2}T_{\mathrm{win}}.$ (58)
This equation gains clarity when we define $T_{k}=1/f_{k}$ as the period of the spectral component with frequency $f_{k}$, becoming
$\nu\ll 2\pi|f_{k}||\tilde{\varphi}(2\pi f_{k})|^{2}\frac{T_{\mathrm{win}}}{T_{k}}.$ (59)
The conditions stated by Eqs. (58) and (59) are independent of the sampling time $T$ and the number of samples $n$ (provided that $n\gg 1$). Specifically, Eq. (59) illustrates that the maximum tolerable linewidth is directly proportional to the modulation frequency, the amplitude square of the phase modulation and the ratio $T_{\mathrm{win}}/T_{k}$, which is the number of temporal periods contained within the time window of the Fourier transform $T_{\mathrm{win}}=nT$. Therefore, widening the time window of the short-time Fourier transform $T_{\mathrm{win}}$ can alleviate the requirements on the laser linewidth compared to the detection of transient phase deviations discussed earlier in this section. However, this can only be achieved by sacrificing temporal resolution, which is determined by the amplitude of the time window $T_{\mathrm{win}}$. Consequently, the necessity of using ultra-stable lasers for detecting low intensity sub-hertz signals persists even with approaches based on the short-time Fourier transform.
The independence of Eqs. (58) and (59) on the sampling time $T$ and the number of samples $n$ suggests that these equations are also valid in the continuous limit where $n$ tends to infinity and $T$ to zero, with their product $T_{\mathrm{win}}=nT$ finite. The derivation of Eqs. (58) and (59) in the continuous case is presented in appendix A.
@@ V Polarization-based sensing
$\langle\tau^{2}\rangle=\frac{1}{\omega_{0}^{2}}\,2L_{\mathrm{f}}\langle\beta^{2}\rangle z,$ (74)
@@
$\langle\Delta\vec{s}(z,t)\cdot\Delta\vec{s}(z,t^{\prime})\rangle=\frac{4}{3}\langle\beta^{2}\rangle L_{\mathrm{f}}\int_{0}^{z}\,4\,\alpha(z^{\prime},t)\alpha(z^{\prime},t^{\prime})\,{\operatorname{d}}z^{\prime},$ (78)
@@
$\langle\Delta\vec{s}(z,t)\cdot\Delta\vec{s}(z,t^{\prime})\rangle=\frac{\pi}{4}\left(\frac{2\pi c}{\lambda}\right)^{2}\kappa^{2}\int_{0}^{z}\,4\,\alpha(z^{\prime},t)\alpha(z^{\prime},t^{\prime})\,{\operatorname{d}}z^{\prime}.$ (79)
@@ VI Comparison between phase and polarization accumulation
[…]
@@ VIII Localization with polarization
Assume that $\mathbf{U}_{\mathrm{f}}$ is the Jones matrix that describes the evolution of the polarization of a single mode fiber up to a given distance $z$ and $\Delta\mathbf{U}$ the unitary matrix describing the evolution in a section that goes from $z$ to $z+\Delta z$ along the same fiber, which we will refer to as the section of interest in the following. Then, the backscattered field is either rerouted through a different fiber with Jones matrix $\mathbf{U}_{\mathrm{b}}$ or transmitted back through the same fiber, in which case $\mathbf{U}_{\mathrm{b}}=\mathbf{U}_{\mathrm{f}}$. The Jones matrix $\mathbf{U}_{\mathrm{rt}}(1)$ describing the round-trip propagation from 0 to $z$ and back, and the Jones matrix $\mathbf{U}_{\mathrm{rt}}(2)$ describing the round-trip propagation from 0 to $z+\Delta z$ and back, areGaltarossa et al. (2008)
@@ IX Environmental sensing using the state of polarization
$\hskip-99.58464pt\left\langle\left[\mathbf{R}_{\omega_{2}}^{-1}(z^{\prime})\vec{\beta}(z^{\prime})\times\vec{s}_{0}\right]\cdot\left[\mathbf{R}_{\omega_{1}}^{-1}(z^{\prime\prime})\vec{\beta}(z^{\prime\prime})\times\vec{s}_{0}\right]\right\rangle.$
The result of the average is (see appendix B for the detailed derivation)
@@
$\hskip-108.12054pt\frac{\pi}{4}\left(\frac{2\pi c}{\lambda}\right)^{2}\kappa^{2}\int_{0}^{z}{\operatorname{d}}z^{\prime}\epsilon(z^{\prime},t^{\prime})\epsilon(z^{\prime},t)F(z^{\prime}),$
@@
[…]
@@
$\hskip-108.12054pt\frac{\pi}{4}\left(\frac{2\pi c}{\lambda}\right)^{2}\xi^{2}\kappa^{2}F(z_{p})\int_{0}^{z}{\operatorname{d}}z^{\prime}\epsilon(z^{\prime},t^{\prime})\epsilon(z^{\prime},t),$
@@
Figure 8: Autocorrelation of $\Delta s_{1}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 552.35 s, from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake 23 June 2020 15:29:05 plotted from 200-200 s to $200$ s (left panel), and a zoom from 50-50 s to $50$ s (right panel).
Figure 9: Autocorrelation of $\Delta s_{2}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 552.35 s, from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake on 23 June 2020, 15:29:05 plotted from 200-200 s to $200$ s (left panel), and a zoom from 50-50 s to $50$ s (right panel).
@@
Figure 16: Autocorrelation of $\Delta s_{1}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 519.9 s (red), from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake, 3 June 2020 07:35:33 plotted from 200-200 s to $200$ s (left panel), and a zoom from 50-50 s to $50$ s (right panel).
Figure 17: Autocorrelation of $\Delta s_{2}$ for channel 1 (blue) with overlapped cross-correlation between channel 1 and channel 2 delayed by 519.9 s (red), from UTC time $t_{0}-1000$ s to $t_{0}+2000$ s where $t_{0}$ is the UTC time of the earthquake on 3 June 2020, 07:35:33 plotted from 200-200 s to $200$ s (left panel), and a zoom from 50-50 s to $50$ s (right panel).
@@
These observations suggest that, in principle, within a single-pass configuration of an operational transmission system, cross-correlation between two closely spaced channels, whose spacing can be optimized for maximum accuracy, may permit the localization of the position on the link where an earthquake occurs. The optimization involves the choice of a frequency spacing maximizing the sensitivity of $F(z_{p})$ on a $z_{p}$ ranging from 0 to the link span $z$. A good recipe may be setting $\Delta z\simeq z$.
@@ X Conclusions
Appendix A Continuous limit of Eqs. (58) and (59)
In this appendix, we derive the continuous limit of Eqs. (58) and (59). After multiplying and dividing by the sampling time $T$, Eqs. (51) and (52) become
$\tilde{N}(\Omega_{k})=\frac{1}{nT}\,\sum_{h=0}^{n-1}N(hT)\exp\left(i\Omega_{k}hT\right)T,$ (129)
$\tilde{\varphi}(\Omega_{k})=\frac{1}{nT}\,\sum_{h=0}^{n-1}\varphi(hT)\exp\left(i\Omega_{k}hT\right)T.$ (130)
Letting $T$ tend to zero and $n$ to infinity while maintaining $nT=T_{\mathrm{win}}$ finite, $T$ becomes ${\operatorname{d}}t$, $hT$ in the exponents becomes a continuous time $t$ and the sums transform into integrals. Thus, we arrive at the continuous limit of these expressions as
$\tilde{N}_{0}(\Omega)=\frac{1}{T_{\mathrm{win}}}\int_{0}^{T_{\mathrm{win}}}\exp(i\Omega t)N_{0}(t){\operatorname{d}}t,$ (131)
and
$\tilde{\varphi}(\Omega)=\frac{1}{T_{\mathrm{win}}}\int_{0}^{T_{\mathrm{win}}}\exp(i\Omega t)\varphi(t){\operatorname{d}}t.$ (132)
Strictly speaking, Eq. (53) dictates that $\Omega$ takes on the discrete values $\Omega=2\pi k/(nT)=2\pi k/T_{\mathrm{win}}$, with $k\in\mathbb{Z}$. We removed the dependence on the integer $k$ because we assume that $\Omega$ is analytically continued over the entire real axis. Again, like $1/n$ in the discrete case, the normalization factor $1/T_{\mathrm{win}}$ ensures that the peak amplitude of the Fourier transform of a sinusoidal modulation is independent of the time window $T_{\mathrm{win}}$. In the continuous limit, the laser phase noise is negligible over the signal if
$|\tilde{N}_{0}(\Omega)|^{2}\ll|\Omega|^{2}|\tilde{\varphi}(\Omega)|^{2}.$ (133)
Inserting into $|\tilde{N}_{0}(\Omega)|^{2}$ the expression (131), averaging the result and using Eq. (47) produces the equality
$\langle|\tilde{N}_{0}(\Omega)|^{2}\rangle=\frac{2\pi\nu}{T_{\mathrm{win}}}.$ (134)
Inserting this expression into Eq. (133) and using the definition $f=\Omega/(2\pi)$ yields
$\nu\ll 2\pi|f|^{2}|\tilde{\varphi}(2\pi f)|^{2}T_{\mathrm{win}},$ (135)
which is the continuous limit of Eq. (58). Using now the definition $f=1/T_{f}$, we obtain the continuous limit of Eq. (59), namely
$\nu\ll 2\pi|f||\tilde{\varphi}(2\pi f)|^{2}\frac{T_{\mathrm{win}}}{T_{f}}.$ (136)
Appendix B Derivation of Eq. (123)
In this appendix, we detail the derivation of Eq. (123). Equation (122) can be rewritten as